Use the separable method to solve the differential equation fracdydx=xy
1
Expert's answer
2019-10-14T11:53:59-0400

Solution: The differential equation using separable method:

"\frac{dy}{dx} = xy"

The equation can be written as follows:


"\frac{dy}{y} = x dx"

We can now integrate both sides:

"\int \frac{dy}{y} = \int x dx"


"\ln y = \frac{1}{2}x^2 + C_1"

"\ln y = \ln e^{\frac{1}{2}x^2 + C_1}"

"y = e^{\frac{1}{2}x^2 + C_1} = e^{\frac{1}{2}x^2} e^{C_1} = C e^{\frac{1}{2}x^2},"


where "C_1" and "C = e^{C_1}" are constants.


Answer:

"y = C e^{\frac{1}{2}x^2}"